OpenStudy (anonymous):

Solve this system of equations using the addition method. x-y=6 -3y=x+22

OpenStudy (anonymous):

ik this aint 2pac

OpenStudy (anonymous):

thats a fail

OpenStudy (anonymous):

\[x-y=6 and -x-3y=22\]

OpenStudy (anonymous):

if you add them you get: \[x-x-y-3y=28\]

OpenStudy (anonymous):

x's are cancelled out and you get: \[-4y=28 and y=\frac{ 28 }{ -4 }=-7\]

OpenStudy (anonymous):

and x=-1

OpenStudy (texaschic101):

-3y = x + 22 (change this to match the other equation) -x - 3y = 22 x - y = 6 --------------add 0 - 4y = 28 -4y = 28 y = - 28/4 = - 7 now sub -7 in for y in either of the equations x - y = 6 x - (-7) = 6 x + 7 = 6 x = 6 - 7 x = - 1 now check it... -3y = x + 22 -3(-7) = -1 + 22 21 = 21 (correct) ANSWER : x = -1, y = -7 or (-1,-7)